OpenStudy (anonymous):

K im having trouble using the integral to find the surface are of y=sqrt(1+4x), 1

OpenStudy (anonymous):

again :)...you're certainly in hell !

OpenStudy (anonymous):

ya, im taking cal 2 over the summer class started on monday first test 2morrow

OpenStudy (anonymous):

\[\int\limits_{1}^{5}f(x)\sqrt{1+ f \prime^2(x) }dx\]

OpenStudy (anonymous):

im pretty lost on this one i got (dy/dx)^2 = 4/(1+4x)

OpenStudy (anonymous):

and that might be my problem

OpenStudy (anonymous):

good luck then ! I've forgotten 2 Pi !

OpenStudy (anonymous):

thank you and ya I know the formulas lol its using them that gets me

OpenStudy (anonymous):

ya if you can just give me a shove into the integral part i think i can take it from there

OpenStudy (anonymous):

yeah..\[\int\limits_{1}^{5}\sqrt{1+4x}*\sqrt{1+4/(1+4x)}dx=\int\limits_{1}^{5}\sqrt{1+4x+4}dx\] you can finish it now ?

OpenStudy (anonymous):

well I see what you did kind of.. but how exactly did you do that? cross multiply?!?!

OpenStudy (anonymous):

did you foil the nominators and that canceled the denominator. haha mind blown

OpenStudy (anonymous):

\[\sqrt{a}\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

and the one on the left is the same thing as sqrt(1+4x)/1 right

OpenStudy (anonymous):

sorry ; I have problem with my connection ! I couldn't post quickly ! aaargh..it takes minutes !!!!

OpenStudy (anonymous):

give me the final answer ! so I can check !

OpenStudy (anonymous):

so Im still kind of confused \[\sqrt{a}\times \sqrt{b/1+4x} = \sqrt{(a(b))/1+4x}\]

OpenStudy (anonymous):

kk! step by step !

OpenStudy (anonymous):

ya, sorry just this part is throwing me for a loop

OpenStudy (anonymous):

\[\sqrt{1+4x}\sqrt{1+4/(4+x)}=\sqrt{(1+4x)1+(1+4x)4/(1+4x)} \]

OpenStudy (anonymous):

(1+4x)4/(1+4x)=4

OpenStudy (anonymous):

no problem ! I'm here for help !

OpenStudy (anonymous):

ok so look at the first message you posted after my sorry message lol the 4 after (1+4x)*4* is that suppose to be there?

OpenStudy (anonymous):

k never mind i got it haha sorry

OpenStudy (anonymous):

so the final answer is.....?!

OpenStudy (anonymous):

k so i got pi/3(5^(3/2) -1)

OpenStudy (anonymous):

does that look good?

OpenStudy (anonymous):

hang on i forgot to sub back in

OpenStudy (anonymous):

sweet! i got 98pi/3 and thats the answer!!!

OpenStudy (anonymous):

no :( !I checked it 5 times...

OpenStudy (anonymous):

yeaaah now right !

OpenStudy (anonymous):

no to 98pi/3

OpenStudy (anonymous):

oh lol ok

OpenStudy (anonymous):

sweet well be on the look out im bound to post a couple more

OpenStudy (anonymous):

w8 w8 !

OpenStudy (anonymous):

actually while i got you the derivative of y=sin(pi(x)) is (picos(pi(x))) right

OpenStudy (anonymous):

seems to be smthing wrong ....Just send me the link kk...

OpenStudy (anonymous):

to my question?

OpenStudy (anonymous):

no your answer !

OpenStudy (anonymous):

and yes...pi cos(pi(x)....Aarghh connection makes me mad !

OpenStudy (anonymous):

ok so how would I square that? would it be pi cos^2(pi(x))

OpenStudy (anonymous):

sorryy !!

OpenStudy (anonymous):

its kewl

OpenStudy (anonymous):

anything els! and you were right 98/3 pi

OpenStudy (anonymous):

right ya how would I square the pi cos(pi(x)) to use it in the same formula? would it just be pi cos^2(pi(x))

OpenStudy (anonymous):

pi^2 no ?!

OpenStudy (anonymous):

so it would be pi^2 cos^2(pi(x))

OpenStudy (anonymous):

yeaaah !

OpenStudy (anonymous):

(ab)^2=a^2b^2

OpenStudy (anonymous):

ok so for this problem its the same thing just with this function y=sin(pi(x)) I've got this

OpenStudy (anonymous):

\[\int\limits_{0}^{1} 2\pi \sin(\pi(x))\]

OpenStudy (anonymous):

oops one sec

OpenStudy (anonymous):

\[\int\limits_{0}^{1} 2\pi \sin(\pi(x))\sqrt{1+(\pi^2 \cos^2(\pi(x))}\]

OpenStudy (anonymous):

k there we go

OpenStudy (anonymous):

yeah ! what do you suggest ?

OpenStudy (anonymous):

really I have no Idea it looks super messy. all I can think of is u substitution put it looks like it will need a substitution within a substitution. I'm hoping not to have to mess with that

OpenStudy (anonymous):

this one is not easy !

OpenStudy (anonymous):

trig is always a problem!

OpenStudy (anonymous):

are you figuring it out man?

OpenStudy (anonymous):

I have an Idea..but I'm not sure yet....cos(x)=u !

OpenStudy (anonymous):

variabl change !

OpenStudy (anonymous):

cos(pix) =u sorry !

OpenStudy (anonymous):

cos(pix)=u It really works ; but I'll give you a better solution ! let G be \[G=\int\limits_{}^{}\sqrt{1+\pi^2u^2}du\] so ! obviously \[G \prime(u)=\sqrt{1+\pi^2u^2} \] then,\[S=2\pi \int\limits_{0}^{1}\sin(\pi x)G \prime(\cos(\pi x))dx=-2\int\limits_{0}^{1}(\cos(\pi x))'G'(\cos(\pi x)dx\] wich equals \[S=-2\left[ \left[ G(\cos(\pi x)\right] \right]\] from 0 to 1 so let's find G using integration by parts \[G=\int\limits_{}^{}(u)'\sqrt{1+\pi^2u^2}du=u \sqrt{1+\pi^2u^2}- \int\limits_{}^{}\pi^2u^2/(\sqrt{1+\pi^2u^2})\] a little tricky right here ! to calculate the second part we add 0=+1-1 \[\int\limits\limits_{}^{}\ (\pi^2u^2+1-1)/(\sqrt{1+\pi^2u^2})du=\int\limits\limits_{}^{}\ (\pi^2u^2+1)/(\sqrt{1+\pi^2u^2})du-\int\limits\limits_{}^{}\ 1/(\sqrt{1+\pi^2u^2})du\] let's begin with the second one ! \[\int\limits_{}^{}1/\sqrt{1+\pi^2u^2}du=1/\pi(\ln \left| u+\sqrt{u^2+\pi^-2} \right|)\] now the firt one !\[\int\limits\limits\limits_{}^{}\ (\pi^2u^2+1)/(\sqrt{1+\pi^2u^2})du =\int\limits_{}^{}\sqrt{1+\pi^2u^2}du=G\] let now put those back to * * \[G=u \sqrt{1+\pi^2u^2}-G-1/\pi(\ln \left| u+\sqrt{u^2+\pi^-2} \right|)\\] \[2G=u \sqrt{1+\pi^2u^2}-1/\pi(\ln \left| u+\sqrt{u^2+\pi^-2} \right|)\ \]

OpenStudy (anonymous):

so \[S=2\pi(G(1)-G(0))\] You owe me a beer ! bad girl !